March 26, 2007

END OF HMT-West 2007

Active field operations for HMT-West 2007 have ended. Unattended sensors will continue to operate, although if they malfunction, they will not be repaired. The plan at present is to remove much of the remaining ESRL instrumentation during April and May. However, some surface sites, particularly those with soil moisture sensors, and the Truckee profiler, will remain in place long term.

Many thanks to the forecasters, engineers, instrument operators and coordinators, chief scientists, program managers and other key individuals that made HMT-West 2007 a success.

Dave Kingsmill

March 25, 2007

Project Status - 25 March 2007

Weather Status: Little change in outlook for precipitation for rain sampling. Rainfall event will start early Monday on the coast and lmid day in the mountains. Peak of event looks to be Monday evening ending by early Tuesday. Showers are possiibile in the cold core period Tuesday afternoon.

System Status: Coastal sampling from Sunday 2130 GMT to Tuesday 2130 GMT, and mountain sampling from Monday 0030GMT to Wed 0030GMT should capture all precipitation.

Operations Status: This is the last operational blog. Thanks to the stalwart few who held on to the end.

John McGinley, GSD

March 24, 2007

Project Status - 24 Mar 2007

Weather Status: Only issue today is when to turn on the rainfall samplers. IPW band still moving down coast but at a slower rate. Precip band moving with front should move into the area Sunday afternoon on the coast and later in the ARB. Cazadero and Bodega Bay look to be about 8 hrs ahead of Alta. Trough should bring moderate amounts of rain to the area.

System Status: Plan is to start the 48-hour collection period at 2130GMT at the coastal stations and 0030GMT Monday at Alta.

John McGinley, GSD

March 23, 2007

Project Status - 23 Mar 2007 - update

Owing to receiver problems at Slough House no soundings will be possible, so all sonde activities will be terminated for this IOP. The IOP will go on for modeling and precipitation collection so there will be a call tomorrow at 1830 GMT. Thanks to the efforts of all the folks at Reno and Slough House for their sonde
support over the HMT exercise.

McGinley

Project Status - 23 Mar 2007

Weather Status: Storm on track for a combination of events: First, an IPW plume is moving down the coast. Observational evidence shows this to be a well defined feature. Models bringing this into the ARB region after 18GMT Sunday, yet forecast no precipitation of any consequence. Second, trough and front with an intense period of warm advection rain begins Monday. Third, cold frontal passage looks to be late Monday/early Tuesday. Total precip amounts look to be in the 2-3 inch category with well defined warm advection precipitation. Thus we are calling an IOP for this event begining at 16GMT Sunday with sonde releases an hour earlier at both Slough House and Reno, and then continuing every four hours. By starting sondes early we can document the establishment of the moist plume, moistening of the PBL, and try to determine if the plume is actually stronger than predicted, and assess why models are not creating any rain in the warm advection on Sunday. Sonde schedules will be adjusted as necesssary at tomorrow's call.

System Status: for planning, sondes at Reno and Slough House centered on Sunday 16, 20; Monday 00, 04, 08 12 16 20; Tuesday 00, 04, 08GMT. Local time releases at Sunday 8am, 12pm, 4pm, 8pm; Monday 12am, 4am, 8am, 12pm 4pm, 8pm; Tuesday 12am....last one.

Rainfall sampling will be done during this event.

John McGinley, GSD

March 22, 2007

Project Status - 22 Mar 2007

Weather Status: Storm for Mon-Tues looking stronger in latest model runs. Precip amounts for the ARB are in the 1-3 inch range. Models indicate significant precipitation commencing early Monday with most intense period Monday evening. Moisture plume now pointed at the Pacific NW will drop south with the approach of the trof and be aimed at the ARB as early as Sunday afternoon. Models are indicating showers begining at this time. So....there are improved prospects for an IOP . One observation strategy would be to document the moistening of the boundary layer with time as the IPW plume and troph merge. This suggests sondes as early as Sunday. A decision on this will be made on Friday. So to summarize: possible IOP Mon-Tues with an outside chance of a Sunday start.

System Status: Sonde inventory shows numbers enough to cover this potential 48-hour event.

Operations Status: With the call today the HMT Experimental Period is formally over, however telecons to discuss this next potential IOP will continue daily at 1830GMT for those interested. The Status and Forecast blogs will also continue through the coming event. Operations are likely to shut down late Tuesday. Thanks to all who participated in the daily calls and the readers of the blogs.

John McGinley, GSD

March 21, 2007

Project Status - 21 Mar 2007

Weather Status: To wrap up the weak event yesterday, the trough did produce some scattered afternoon showers and thunderstorms that produced some moderate precipitation in the Central Valley. Blue Canyon had a few hundredths out of an afternoon shower. Focus for today is a system due to impact the ARB on Mon-Tues (Mar 26-27). If systems maintains its forecast strength HMT will be extended to capture the event. Flow by the weekend is predicted to be more zonal with a weak but well defined moisture plume and good wind speeds at 700mb. Most precipitation by the GFS appears to be frontal. The storm will have abundant cold air so snow levels may be low for March. Liquid amounts look like 1-2 inches, but this appears to be a downward trend relative to earlier progs. Timing is pretty consistent within the GFS ensemble, but the ECMWF is slower by 24 hrs. Timing on event looks like first raindrops at 18GMT on Mon (Mar 26) and ending Tues at 12GMT. So another brief event. Storm is still a monitoring target so we'll update the prospects tomorrow.

System Status: ALPS timing tests were completed for the RFC and RNO, still some problems with SAC. MRY will be done tomorrow.

John McGinley, GSD

March 20, 2007

Project Status - 20 Mar 2007

Weather Status: Storm over ARB almost a complete bust as wave basically split with energy going north and south of the experimental area. As of 20GMT Blue Canyon had no observed precipitation, despite the fact that the NAM last night was going for near an inch. Model ensembles never quit forecasting precipitation as well with amounts up to 1.5 inches even as late as the 20/06GMT runs. This mornings runs hinting at shower activity this afternoon beneath the cold core. NWS also expecting showers, so there may not be a complete shutout as the afternoon progresses. Case will be saved to study a bust situation. Beyond the end of the exercise the storm for next Mon-Tues (26-27 March) looks to be on an upward trend in the GFS and other global models. So outlook is for a possible event Monday-Tuesday. This will extend the exercise 4-5 days for ESRL staff.

System Status: Lack of rain at Slough House precluded the receiver test. All models up and running but not forecasting well.

Operations beyond 22 March: Telecons will be conducted with only ESRL personnel. Chief Scientist Fri-Tuesday: McGinley; Forecaster Fri-Sun: Szoke; Forecaster Mon-Tue: Schultz. Call participants will be all ESRL HMT staff: Kingsmill, Martner, Schneider, Neiman, White, Splitt, as available.

John McGinley, GSD

March 19, 2007

Project Status - 19 Mar 2007

Weather Status: Storm approaching coast is on schedule for a Tuesday morning arrival in the ARB, however precipitation amounts are marginal at best. Models keeping precip in the ARB in the .5-0.75 range over a very short duration (12-15 hrs), so not calling an IOP still looks like a good decision at this time. Ensemble models showing a bit more precip than large scale models with amounts up to 2 inches 100km NW of the ARB. In the ARB amounts are below an inch. NWS is in agreement with 0.5-0.75 forecast for the ARB. Further out, a very strong jet in the mid Pacific coupled to increasing convection in the area near 140E will begin to change things in the eastern Pacific. Long range models and MJO/Rossby wave propagation diagnostics indicate troughy conditions developing over the coast over the next two weeks. This, of course is beyond the sheduled end of the HMT. Consensus was that two storms (26-27 and 29-30 March) should be monitored for possible data collection. So...no IOP on current wave. Extension of project into next week is possible.

System Status: All local models back in action so a workstation comms timing test will be scheduled for Tuesday.

Operations Proposal: Daily telecons will end on the 22nd. ESRL personnel (only) will have a daily local consultation 23 through 30 March to determine if soundings should be launched for the two storms likely in the next week (26-27 March and 29-30 March). NWS and Slough House will be notified 24-48 hrs in advance for supplemental sondes. NWS will be consulted on forecast details as needed. ESRL personnel will will serve as volunteer forecasters through the extension period. Please give this proposal some condideration. We will discuss it on Thursday.

John McGinley, GSD